# FREE-Merging: Fourier Transform for Model Merging with Lightweight Experts

### v1

@@ 6 Conclusion
In this paper, we propose FR-Merging and FREE-Merging to efficiently merge models, leveraging existing models to create a multi-task model. FR-Merging builds a backbone without training by high-pass filtering to remove information that hinders generalization, while FREE-Merging employs lightweight experts to offset information loss during merging. Our method balances training, storage, inference costs, and performance, showing high practical value.
@@ 9.3 Details
10 More Discussions and Future Directions
10.1 Role of Task Experts
@@
10.2 Task Experts Construction Analysis
@@
This section discusses the rationale behind using the parameters with the largest changes during the fine-tuning to approximate task-specific information introduced by fine-tuning, thereby constructing task experts.By comparing L1 distances of model parameters in the frequency domain, with smaller distances indicating better approximation, we assess the effectiveness of different methods.
@@
10.3 Filtering Frequency Analysis
@@
10.4 Future Works
In this section, we outline directions for our future work. First, we aim to extend our proposed model merging approach to fine-tuned models with different architectures, enabling broader applicability. Additionally, we plan to further investigate the differences between PEFT and full fine-tuning and design customized model merging methods tailored to PEFT, enabling effective utilization and deployment of existing PEFT results.
11 Additional Experimental Results
11.1 Merging Vision Models
@@
11.2 Merging Language Models

### v2

@@ 7 Conclusion
In this paper, we propose FR-Merging and FREE-Merging to efficiently merge existing models to create a multi-task model. FR-Merging constructs a backbone without training by high-pass filtering to remove information that dominates task interference. Meanwhile, FREE-Merging employs lightweight experts to offset information loss during merging. Our method balances training, storage, inference costs, and performance, showing high practical value.
@@ C.3 Details
Figure 6: The relationship between the merging performance of ViT-B-32 and the variance of frequency domain amplitude power. Normalized accuracy is the ratio of merged model performance to fine-tuned model performance on the same task.
Appendix D More Discussions and Future Directions
In this section, we first provide a more detailed discussion of task interference in the frequency domain in D.1. Next, we present the experimental results of FR-Merging on language models in D.2 and D.3. Additionally, we discuss the effectiveness and construction of the expert module in D.4 and D.5, as well as the role of the router in D.6. We also explore the selection of the cutoff frequency in D.7 and conclude with a discussion of future works in D.8.
D.1 Task Interference in Frequency Domain
In this section, we further discuss the task interference presented in Sec. 4 and provide additional visualization results.
The first point to address is why we consider the difference in low-frequency amplitude in the frequency domain as the cause of task interference. The amplitude in the frequency domain represents the strength of a signal. During the merging process, if one signal is strong while the other is weak, the weaker signal is likely to be overshadowed by the stronger one, regardless of their directional relationship. This manifests in the merged model as the near-complete loss of performance from one of the fine-tuned models. Therefore, a significant difference in the low-frequency amplitude will indeed lead to task interference during merging.
In Sec. 4, we quantify interference using the mean variance, observing a strong negative correlation with merged model performance on ViT-B/32. Here, normalized performance refers to the ratio of the performance of merged model on a task to that of the fine-tuned model on the same task. This negative correlation serves as strong evidence of task interference in the frequency domain.
Next, we verify whether this negative correlation also holds for language models. We choose Flan-T5 [6] as the base model and conduct experiments on eight GLUE datasets. As shown in Fig. 6, the negative correlation is observed in language models, demonstrating that task interference in the frequency domain is a general phenomenon.
Figure 7: The parameter frequency domain distribution of ViT-B-32 fine-tuning results on eight image tasks.
Finally, we examine whether low-frequency task interference is widespread across different models. In Fig. 7 and Fig. 8, we visualize the frequency distribution of fine-tuned parameters for ViT-B/32 and Flan-T5 [6] on their respective tasks. The results show that task interference in the low-frequency region is consistently present across various models. Therefore, our hypothesis is well-supported. Meanwhile, we observe an interesting phenomenon: in some layers, frequency-domain amplitude clustering occurs, meaning that certain models have very similar amplitude distributions. This may be related to the training path during the finetuning process, and we will further investigate this phenomenon in future research.
Figure 8: The parameter frequency domain distribution of Flan-T5 fine-tuning results on eight GLUE tasks.
Figure 9: a) Merge by preserving the top 20% of parameters; b) Merge using the DARE method; c) FR-Merging. a-c complete on ViT/B-32 and test on 8 visual tasks. d) High-pass filtering improves performance over the unfiltered version on RoBERTa and test on GLUE.
D.2 Expanded Loss Landscape Analysis
In this section, we attempt to analyze why our proposed FR-Merging outperforms existing approaches from the perspective of the loss landscape. A loss landscape comparison with two popular sparse methods, Top-K retention and DARE, is shown in Fig. 9 (a-c). Our proposed FR-Merging method brings the two fine-tuned models closer in the loss landscape for the same task, making their linear combination fall closer to the loss basin of the target task, thereby achieving better performance. This aligns with our analysis from the perspective of task interference and is also consistent with the performance improvements in experiments. Therefore, our proposed FR-Merging method offers significant advantages over current approaches.
D.3 Fourier Transform on Language Models
In Sec. 5.2, we analyze the impact of high-pass filtering on the generalization ability of fine-tuned models and their performance on corresponding tasks using ViT-B/32. In this part, we extend the analysis to language models, further demonstrating the effectiveness of high-pass filtering. Fig. 9(d) shows the generalization test of RoBERTa on the GLUE datasets, where filtering out low-frequency information also enhances the generalization ability of fine-tuned language models. This improvement is also reflected in the merging performance improvement for language models. However, compared to the effect on language models, high-pass filtering has a more pronounced impact on ViT. This may be attributed to the nature of image data, which inherently possesses certain high-frequency and low-frequency properties. Nevertheless, in summary, FR-Merging, based on high-pass filtering, is a generalizable, efficient, and high-performing model merging approach.
D.4 Role of Task Experts
@@
D.5 Task Experts Construction Analysis
@@
In this section, we analyze the construction of experts from two aspects: the parameter selection and rescaling.
Top-K Selection. We discuss the rationale behind using the parameters with the largest changes during the fine-tuning to approximate task-specific information introduced by fine-tuning, thereby constructing task experts.By comparing L1 distances of model parameters in the frequency domain, with smaller distances indicating better approximation, we assess the effectiveness of different methods.
@@
ViT-B/32 ViT-B/16 ViT-L/14
w/o scaling 0.5314 0.5323 0.6090
w/ scaling 0.8484 0.8486 0.8895
Table 12: Ablation of scaling for experts on 8 visual tasks.
Rescaling. We then discuss the necessity of rescaling. Since we retain only a small portion of the parameters during expert extraction, it alters the mean and variance of parameters, similar to dropout. Our scaling method considers both the mean and the backbone to preserve the similarity between the expert and original model outputs. Table 12 shows the CKA similarity between expert outputs and original outputs before and after scaling, with higher values indicating better expert performance. Scaling can be found to greatly improve performance compared to not applying it.
D.6 Role of Router
In this section, we discuss why a router is necessary for dynamically assigning input data to the appropriate expert. First, let’s review how the router works: when a new data arrives, the router determines which existing expert is best suited to handle it. For example, if we have experts for mathematics and physics, and a math problem appears, the router will automatically route it to the mathematics expert. The mathematics expert will then collaborate with the merged backbone to solve the problem. Manually assigning each input sample to the correct task is impractical, and in real-world applications, we cannot expect users to know in advance which expert is best suited for each piece of data. Therefore, having a router to automate this process is both crucial and aligned with real-world usage scenarios.
D.7 Filtering Frequency Hyperparameter Analysis
@@
D.8 Future Works
In this section, we outline directions for our future work. First, we aim to extend our merging methods to fine-tuned models with different architectures, enabling broader applicability. Additionally, we plan to further investigate the differences between PEFT and full fine-tuning and design customized merging methods tailored to PEFT, enabling effective utilization and deployment of existing PEFT results.
Appendix E Additional Experimental Results
E.1 Merging Vision Models
@@
E.2 Merging Language Models